ROTATIONAL SPECTRA OF PHENYLALANINE, TIROSINE AND TRYPTOPHAN

S. MATA, C. PEREZ, M. E. SANZ, \underline{S. BLANCO}, J. C. LÓPEZ, and J. L. ALONSO, Grupo de Espectroscopía Molecular (GEM), Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47005 Valladolid, Spain.

The rotational spectra of the aromatic natural amino acids phenylalanine, tyrosine and tryptophan have been investigated by Laser Ablation Molecular Beam Fourier transform Microwave Spectroscopy LA-MB-FTMW. The spectra of two rotamers of phenylalanine have been detected in the supersonic expansion. Both forms are stabilized by a chain of intramolecular hydrogen bonds O-H⋯N-H⋯π, being the carboxylic group incis configuration. One conformer of tyrosine, which only differs from phenylalanine in a -OH group inpara position, has been also characterized. Preliminary results on the rotational spectrum of tryptophan are presented.